# LLMs Are Biased Towards Output Formats! Systematically Evaluating and Mitigating Output Format Bias of LLMs

### v1

@@ Abstract
We present the first systematic evaluation examining format bias in performance of large language models (LLMs).  [^1] Our codes and data will be made publicly available at link. Our approach distinguishes between two categories of an evaluation metric under format constraints to reliably and accurately assess performance: one measures performance when format constraints are adhered to, while the other evaluates performance regardless of constraint adherence. We then define a metric for measuring the format bias of LLMs and establish effective strategies to reduce it. Subsequently, we present our empirical format bias evaluation spanning four commonly used categories—multiple-choice question-answer, wrapping, list, and mapping—covering $15$ widely-used formats. Our evaluation on eight generation tasks uncovers significant format bias across state-of-the-art LLMs. We further discover that improving the format-instruction following capabilities of LLMs across formats potentially reduces format bias. Based on our evaluation findings, we study prompting and fine-tuning with synthesized format data techniques to mitigate format bias. Our methods successfully reduce the variance in ChatGPT’s performance among wrapping formats from $235.33$ to $0.71$ (%$\%^{2}$).
@@ 1 Introduction
    We introduce the first systematic framework to evaluate format performance bias in LLMs.
@@
    A large-scale evaluation spanning $15$ formats, $8$ tasks, and $3$ models revealing substantial LLM performance variance across formats.
@@ 2 Related Works
Large language models (LLMs) have shown remarkable proficiency in formatting outputs to meet human expectations. Such formats include markdown for lists and pointers (Achiam et al., 2023), code blocks (Gur et al., 2022), and integrate tags, or LaTeX for scientific texts (Singh et al., 2023; Wang et al., 2024). Given the rising importance of formatting capabilities in LLMs, recently, format-following benchmarks have been developed for assessing LLMs’ adherence to specified formats (Zhou et al., 2023; Xia et al., 2024; Chen et al., 2024; Macedo et al., 2024; Liu et al., 2024). However, these studies only evaluate format-instruction following capabilities. Our research further assesses LLM performance across different formats, uncovering significant format bias in various tasks and models. We also acknowledge the concurrent work by Tam et al. (2024), which examines the impact of format restrictions on LLM performance. However, unlike our approach, they do not disentangle evaluation metrics under format constraints and only evaluate $3$ structured formats, substantially fewer than our study.
@@ 3.3 Formats for Evaluation
where the output of LLMs is a list of elements. We investigate $4$ formats representing lists: (1) Python list Do et al. (2023); (2) Bullet-point list Liu et al. (2024); (3) List of elements separated by a special character ”[SEP]” Boucher (2023); and (4) List of elements arranged on separate lines Mishra (2023).
@@ 4 General Experimental Setups
For MCQ bias evaluation (Section 5.1), we select two datasets: MMLU (Hendrycks et al., 2021) and BBH (Suzgun et al., 2023). For MMLU, we randomly choose 27 subcategories. For BBH, we select the sports_understanding category following Gupta et al. (2024). For wrapping bias assessment (Section 5.2), in addition to MCQ benchmarks, the following datasets are experimented: GSM8K (Cobbe et al., 2021) for reasoning, FairytaleQA (Xu et al., 2022a) for narrative comprehension, and HotpotQA (Yang et al., 2018a) for multi-hop reasoning. For list bias investigation (Section 5.3), we use SciDocsRR (Muennighoff et al., 2023), a scientific document ranking task as the order list generation task, and SemEval 2017 (Augenstein et al., 2017a), the keyphrase extraction task as the unordered list generation. For mapping bias examination (Section 5.4), we utilize a document-level information extraction task named SciREX (Jain et al., 2020a) by synthesizing three extraction difficulty levels: easy (extracting from $1$ sentence for $1$ category), medium ($3$ sentences, $2$ categories), and hard ($5$ sentences, $4$ categories). For all benchmarks except MCQ, we sample $200$ points for evaluation Bai et al. (2023).
@@
We select both open- and closed-source LLMs for our evaluation: Gemma-7B-it (Team et al., 2024) and Mistral-7B-it-v0.2 (Jiang et al., 2023) for open-source as they are among state-of-the-art open-source LLMs; ChatGPT (gpt-3.5-turbo-0125) for closed-source as this premier chatbot possesses superior instruction-following ability. Our purpose is not to reproduce the models’ performance, but to show the bias.
@@ 5 Format Evaluation Experiments
Overall, we find that: (1) Models show substantial format-following bias across formats for all benchmarks; (2) For all models and datasets, significant performance bias exists across formats; (3) $77.67\%$ of the $EstTrue$ results are reliable, with $16/24$ for MCQ, $169/210$ for wrapping, $35/48$ for list, and $27/36$ for mapping formats. We dive into (2) for every format as it is our main focus, (1, 3) are discussed in detail in Sections C.1, C.2, C.3 and C.4.
@@ Results.
Fig. 2 provides a synopsis of our evaluation results, with numerical values shown in Appx.-Tab. 2. From Fig. 2-left, we observe that Mistral possesses the highest disparity between the two MCQ answer formats, with $58.69\%$ accuracy on average for character and only $4.22\%$ for textual value. Additionally, despite ChatGPT often being regarded as one of the most robust LLMs, it shows a significant performance difference between the two formats ($19.03\%$). Overall, LLMs are heavily biased towards outputting character identifiers. Requiring them to generate the choice’s value causes notable performance drops of $28.76\%$ on average.
@@
Fig. 3 outlines an overview of our evaluation outcomes with results in Appx.-Tab. 6. From Fig. 3-left, we see that Gemma exhibits the highest bias towards different formats with a $BiasF_{o}$ value (Eq. 8 (variance) of $56.33\%^{2}$, while ChatGPT performs the best with only $12.26\%^{2}$. Notably, for “Quoting” and “Parenthesis”, the Gemma follows instructions only about $0-4\%$ yielding nearly zero performance, highlighting its critical weaknesses. Among the 7 formats, “Placeholder” ($35.92\%$) proves to be the most effective wrapping output format, while “Quoting” ($23.74\%$), “Parenthesis” ($28.64\%$) are among those that achieve the lowest performance.
From Fig. 3-right, models exhibit bias across all tasks, with the lowest on GSM8K ($12.97\%^{2}$) possibly because the models were trained on (part of) it, and the highest on BBH ($70.16\%^{2}$), the challenging task without train data. This demonstrates the pervasive presence of wrapping bias in LLMs.
@@
The format token bias of LLMs as explained in Section 5.1 is also our hypothesis. Specifically, we found the low performance of the “Quoting” and ”Parenthesis” because, in generation tasks, models often wrap (via quoting/parenthesizing) not only the final answer, as instructed, but also parts of the context (e.g., “‘The answer is 3.”’), leading to poor F1 scores. Moreover, Gemma completely ignores the above format instructions, resulting in 0% FI scores, which also contribute to the low average estimated F1 scores. These strongly indicate the presence of format token bias.
@@
Fig. 4 displays the key findings of our evaluation across models and datasets with numerical results in Appx.-Tab. 10. From Fig. 4-left, we notice that Mistral exhibits the most bias, with the $BiasF_{o}$ value (Eq. 8) of $353.80\%^{2}$. In contrast, ChatGPT and Gemma show much lower bias, with values of $7.08\%^{2}$ and $1.32\%^{2}$, respectively. Of the four formats, the “Python” and “Newline-separated” formats yield the highest performance, likely due to models trained extensively on code data. Conversely, the “Bullet-point list” format results in the lowest performance, particularly for Mistral, highlighting the inherent bias for such formats.
@@
The performance bias is regardless of the task as plotted in Fig. 4-right, with the highest $BiasF_{o}$ value of $67.07\%^{2}$ on the order list generation task SciDocsRR, and significantly lower ($27.58\%^{2}$) on SemEval2017 task. The high bias in the SciDocsRR task is because Mistral and Gemma mostly failed to perform this task following the “Bullet” and “Special character” list formats while excelling in solving it following the other formats.
@@ 5.4 Experiments on Mapping Format
We examine the performance bias of LLMs on two mapping formats as discussed in Section 3: (1) Python dictionary/JSON; (2) YAML. We preprocess the SciREX task Jain et al. (2020a) as described in Section 4 into three extraction levels: (i) Easy (1 sentence, “Task” category); (2) Medium (3 sentences, “Task, Method”); (3) Hard (5 sentences, “Task, Method, Material, Metric” categories).
@@ Results.
Fig. 5 illustrates a summary of our evaluation with numerical details in Appdx.-Tab. 14. From Fig. 5-left, Gemma is the most biased, with a performance gap of $16.51\%$ between the two formats, followed by Mistral with a $16.07\%$ gap. ChatGPT, however, is relatively robust against format variations, exhibiting a gap of only $4.05\%$. On average, JSON performs significantly better than YAML for mapping, likely because more JSON data is used to train models due to its popularity.
@@ 6 Mitigating Performance Format Bias: Actionable Recommendations
We propose methods as actionable recommendations for mitigating format bias. Generally, three primary streams of techniques have been widely studied and applied to tackle LM biases: (1) Prompting Xu et al. (2024); Macedo et al. (2024); (2) Calibrating Roelofs et al. (2022); Li et al. (2024); and (3) Fine-tuning Schick et al. (2021); Ghaddar et al. (2021). While calibration techniques can only be used for white-box models, prompting and fine-tuning can be applied for both black-box (via API) and white-box ones. Therefore, we explore prompting and fine-tuning techniques to reduce format bias. We target mitigating the format bias of ChatGPT, the strongest model that we benchmarked, on MMLU. We aim to reduce the wrapping bias (Section 5.2) due to resource limits, but our methods can be generalized to any model and format.
@@
As discussed in Section 5.1, LLMs show bias across formats possibly because of the token bias issue, causing LLMs to non-uniformly comprehend the format instructions. To address this, we examine whether demonstrations with formats can reduce such bias, as they are commonly utilized to enhance LLM’s comprehension of the task patterns Xie et al. (2022). Particularly, for each wrapping format in Section 5.2, we select $1$ and $5$ random samples from the auxiliary train data of MMLU and manually format the answers as demonstrations. The results are outlined in Fig. 6. Firstly, incorporating demonstrations typically enhances the FI scores (i) (from 72.99% to 79.19% and 85.31%) of the model, with five demonstrations yielding the most. Secondly, we observe a notable decrease in the $BiasF_{o}$ score (ii) upon supplementing demonstrations. From (i), (ii) and Section 3.2, we conclude integrating demonstrations mitigates format bias.
@@ Acknowledgements
This research is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC2023-010-SGIL) and the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207). DXL and TS are supported by the A*STAR Computing and Information Science (ACIS) scholarship. We thank Goh Yisheng for his contribution in the initial stage of the project. We thank members of WING and Deep Learning Lab, NUS, as well as the anonymous reviewers for the constructive feedback.
@@ Appendix B Theoretical Analysis: Reliability of $EstTrueE$
$\lim_{{FI_{C}}\to 100\%}{(EstTrueE)}=EstTrueE(FI_{C}=100\%)={TrueE}.$
@@ C.1 Multiple-choice Question (MCQ) Discussions
Table 2: MCQ output format results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
MMLU 39.56 26.36 43.56
BBH 55.77 11.46 490.84
Table 3: Average. estimated true accuracy results of MCQ benchmarks across models (Gemma-7B-it, Mistral, ChatGPT) and datasets (MMLU, BBH) with two MCQ formats (Character and Textual Value). This table is the distillation result of averaging all the estimated true performance scores across models (Models) and tasks (Tasks) in Tab. 3.
@@
The FI scores across formats are illustrated in Tab. 4. There is a notable difference between the scores for character-based and textual value-based formats. The score for character format (51.39%) is more than twice as high as that for textual value format (20.46%). Among the models, ChatGPT follows the instructions best with FI score $52.42\%$. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
@@
We investigate the impact of Chain-of-Thought (CoT) Wei et al. (2022) on mitigating format bias. The average of estimated True Accuracy scores across models and tasks for the ZS prompting and ZS-CoT prompting are shown in Tab. 5. ZS prompting achieves a higher score in the character category (51.20%) compared to ZS-CoT prompting (44.13%). Similarly, for textual value format, ZS prompting scores higher (20.83%) than ZS-CoT prompting (16.99%). However, the $BiasF_{o}$ is lower for the ZS-CoT model ($184.14\%^{2}$) compared to the ZS model ($230.58\%^{2}$), indicating that CoT slightly decreases the format bias.
@@
Zero-shot 51.20 20.83 230.58
Zero-shot Chain-of-Thought 44.13 16.99 184.14
@@
From Tab. 2, we see that $16/24$ of the estimated $EstTrue$ results are reliable. The reliability of results in the MCQ output format varies across different models. Gemma-7B-it and Mistral-7B-it show significant unreliability in textual value format, evidenced by numerous red-marked scores due to models not following the format instructions to output correct formats. In contrast, ChatGPT’s results are significantly more reliable in the MMLU and BBH benchmarks ($7/8$), with only one unreliable result in the BBH textual format output.

### v2

@@ Abstract
We present the first systematic evaluation examining format bias in performance of large language models (LLMs). Our approach distinguishes between two categories of an evaluation metric under format constraints to reliably and accurately assess performance: one measures performance when format constraints are adhered to, while the other evaluates performance regardless of constraint adherence. We then define a metric for measuring the format bias of LLMs and establish effective strategies to reduce it. Subsequently, we present our empirical format bias evaluation spanning four commonly used categories—multiple-choice question-answer, wrapping, list, and mapping—covering $15$ widely-used formats. Our evaluation on eight generation tasks uncovers significant format bias across state-of-the-art LLMs. We further discover that improving the format-instruction following capabilities of LLMs across formats potentially reduces format bias. Based on our evaluation findings, we study prompting and fine-tuning with synthesized format data techniques to mitigate format bias. Our methods successfully reduce the variance in ChatGPT’s performance among wrapping formats from $235.33$ to $0.71$ (%$\%^{2}$).
@@ 1 Introduction
    We introduce the first systematic framework [^1] Our codes and data will be made publicly available at link. to evaluate format performance bias in LLMs.
@@
    A large-scale evaluation spanning $15$ formats, $8$ tasks, and $4$ models revealing substantial LLM performance variance across formats.
@@ 2 Related Works
Large language models (LLMs) have shown remarkable proficiency in formatting outputs to meet human expectations. Such formats include markdown for lists and pointers (Achiam et al., 2023), code blocks (Gur et al., 2023), and integrate tags, or LaTeX for scientific texts (Singh et al., 2023; Wang et al., 2024). Given the rising importance of formatting capabilities in LLMs, recently, format-following benchmarks have been developed for assessing LLMs’ adherence to specified formats (Zhou et al., 2023; Xia et al., 2024; Chen et al., 2024; Macedo et al., 2024; Liu et al., 2024). However, these studies only evaluate format-instruction following capabilities. Our research further assesses LLM performance across different formats, uncovering significant format bias in various tasks and models. We also acknowledge the concurrent work by Tam et al. (2024), which examines the impact of format restrictions on LLM performance. However, unlike our approach, they do not disentangle evaluation metrics under format constraints and only evaluate $3$ structured formats, substantially fewer than our study.
@@ 3.3 Formats for Evaluation
where the output of LLMs is a list of elements. We investigate $4$ formats representing lists: (1) Python list Do et al. (2025); (2) Bullet-point list Liu et al. (2024); (3) List of elements separated by a special character “[SEP]” Boucher (2023); and (4) List of elements arranged on separate lines Mishra (2023).
@@ 4 General Experimental Setups
For MCQ bias evaluation (Section 5.1), we select two datasets: MMLU (Hendrycks et al., 2021) and BBH (Suzgun et al., 2023). For MMLU, we randomly choose 27 subcategories. For BBH, we select the sports_understanding category following Gupta et al. (2024). For wrapping bias assessment (Section 5.2), in addition to MCQ benchmarks, the following datasets are experimented: GSM8K (Cobbe et al., 2021) for reasoning, FairytaleQA (Xu et al., 2022) for narrative comprehension, and HotpotQA (Yang et al., 2018) for multi-hop reasoning. For list bias investigation (Section 5.3), we use SciDocsRR (Muennighoff et al., 2023), a scientific document ranking task as the order list generation task, and SemEval 2017 (Augenstein et al., 2017), the keyphrase extraction task as the unordered list generation. For mapping bias examination (Section 5.4), we utilize a document-level information extraction task named SciREX (Jain et al., 2020) by synthesizing three extraction difficulty levels: easy (extracting from $1$ sentence for $1$ category), medium ($3$ sentences, $2$ categories), and hard ($5$ sentences, $4$ categories). For all benchmarks except MCQ, we sample $200$ points for evaluation Bai et al. (2024).
@@
We select both open- and closed-source LLMs for our evaluation: Gemma-7B-it (Team et al., 2024), Mistral-7B-it-v0.2 (Jiang et al., 2023), and Llama-3.1-8B-it (Dubey et al., 2024) for open-source as they are among state-of-the-art open-source LLMs; ChatGPT (gpt-3.5-turbo-0125) for closed-source as this premier chatbot possesses superior instruction-following ability. Our purpose is not to reproduce the models’ performance, but to show the bias.
@@ 5 Format Evaluation Experiments
Overall, we find that: (1) Models show substantial format-following bias across formats for all benchmarks; (2) For all models and datasets, significant performance bias exists across formats; (3) $78.30\%$ of the $EstTrue$ results are reliable (with $70\%$ for MCQ, $82.5\%$ for wrapping, $67.19\%$ for list, and $77.08\%$ for mapping formats) highlighting significant weaknesses of LLMs in following format instructions. We dive into (2) for every format as it is our main focus, (1, 3) are discussed in detail in Sections C.1, C.2, C.3 and C.4.
@@ Results.
Fig. 2 provides a synopsis of our evaluation results, with numerical values shown in Appx.-Tab. 2. From Fig. 2-left, we observe that Mistral possesses the highest disparity between the two MCQ answer formats, with $58.69\%$ accuracy on average for character and only $4.22\%$ for textual value. Additionally, despite ChatGPT often being regarded as one of the most robust LLMs, it shows a significant performance difference between the two formats ($19.03\%$). Overall, LLMs are heavily biased towards outputting character identifiers. Requiring them to generate the choice’s value causes notable performance drops on most models.
@@
Fig. 3 outlines an overview of our evaluation outcomes with results in Appx.-Tab. 6. From Fig. 3-left, we see that Llama exhibits the highest bias towards different formats (with a $BiasF_{o}$ value of $74.86\%^{2}$; see Appx.-Tab. 7), while ChatGPT performs the best. Notably, for “Quoting” and “Parenthesis”, the Gemma follows instructions only about $0-4\%$ yielding nearly zero performance, highlighting its critical weaknesses. Among the 7 formats, “Placeholder” ($37.15\%$) proves to be the most effective wrapping output format, while “Quoting” ($24.58\%$), “Parenthesis” ($28.57\%$) are among those that achieve the lowest performance.
From Fig. 3-right, models exhibit bias across all tasks, with the lowest on MMLU ($16.58\%^{2}$; see Appx.-Tab. 7) possibly because the models already performed relatively well on it, and the highest on BBH ($54.58\%^{2}$), the challenging task without train data. This demonstrates the pervasive presence of wrapping bias in LLMs.
@@
The format token bias of LLMs as explained in Section 5.1 is also our hypothesis. Specifically, we found the low performance of the “Quoting” and “Parenthesis” because, in generation tasks, models often wrap (via quoting/parenthesizing) not only the final answer, as instructed, but also parts of the context (e.g., “‘The answer is 3.”’), leading to poor F1 scores. Moreover, Gemma completely ignores the above format instructions, resulting in 0% FI scores, which also contribute to the low average estimated F1 scores. These strongly indicate the presence of format token bias.
@@
Fig. 4 displays the key findings of our evaluation across models and datasets with numerical results in Appx.-Tab. 10. From Fig. 4-left, we notice that Mistral exhibits the most bias, with the $BiasF_{o}$ value of $353.80\%^{2}$. In contrast, ChatGPT and Gemma show much lower bias, with values of $7.08\%^{2}$ and $1.32\%^{2}$, respectively. Of the four formats, the “Python” and “Newline-separated” formats yield the highest performance, likely due to models trained extensively on code data. Conversely, the “Bullet-point list” format results in the lowest performance, particularly for Mistral, highlighting the inherent bias for such formats.
@@
The performance bias is regardless of the task as plotted in Fig. 4-right, with the highest $BiasF_{o}$ value of $54.12\%^{2}$ on the order list generation task SciDocsRR, and significantly lower ($31.86\%^{2}$) on SemEval2017 task. The high bias in the SciDocsRR task is because Mistral and Gemma mostly failed to perform this task following the “Bullet” and “Special character” list formats while excelling in solving it following the other formats.
@@ 5.4 Experiments on Mapping Format
We examine the performance bias of LLMs on two mapping formats as discussed in Section 3: (1) Python dictionary/JSON; (2) YAML. We preprocess the SciREX task Jain et al. (2020) as described in Section 4 into three extraction levels: (i) Easy (1 sentence, “Task” category); (2) Medium (3 sentences, “Task, Method”); (3) Hard (5 sentences, “Task, Method, Material, Metric” categories).
@@ Results.
Fig. 5 illustrates a summary of our evaluation with numerical details in Appdx.-Tab. 14. From Fig. 5-left, Gemma is the most biased, with a performance gap of $16.51\%$ between the two formats, followed by Mistral with a $16.07\%$ gap. ChatGPT and Llama, however, are relatively robust against format variations. On average, JSON performs significantly better than YAML for mapping, likely because more JSON data is used to train models due to its popularity.
@@ 6 Mitigating Performance Format Bias: Actionable Recommendations
We propose methods as actionable recommendations to mitigate format bias in the performance of LLMs. Generally, three primary streams of techniques have been widely studied and applied to tackle LM biases: (1) Prompting Xu et al. (2024); Macedo et al. (2024); (2) Calibrating Roelofs et al. (2022); Li et al. (2024); and (3) Fine-tuning Schick et al. (2021); Ghaddar et al. (2021). While calibration techniques can only be used for white-box models, prompting and fine-tuning can be applied for both black-box (via API) and white-box ones. Therefore, we explore prompting and fine-tuning techniques to reduce format bias. We target mitigating the format bias of ChatGPT in Fig. 6, the strongest model that we benchmarked, on MMLU. We aim to reduce the wrapping bias (Section 5.2) due to resource limits, but our methods can be generalized to any model and format. We also verify (and confirm) our mitigation strategies on Gemma-2B-it, a medium-size model in Appx.-Tab. 18.
@@
As discussed in Section 5.1, LLMs show bias across formats possibly because of the token bias issue, causing LLMs to non-uniformly comprehend the format instructions. To address this, we examine whether demonstrations with formats can reduce such bias, as they are commonly utilized to enhance LLM’s comprehension of the task patterns Xie et al. (2022). Particularly, for each wrapping format in Section 5.2, we select $1$ and $5$ random samples from the auxiliary train data of MMLU and manually format the answers as demonstrations. The results are outlined in Fig. 6. Firstly, incorporating demonstrations typically enhances the FI scores of the model (from 72.99% to 79.19% and 85.31%) (i), with five demonstrations yielding the most. Secondly, we observe a notable decrease in the $BiasF_{o}$ score (ii) upon supplementing demonstrations. From (i), (ii) and Section 3.2, we conclude integrating demonstrations mitigates format bias.
@@ Acknowledgements
This research project is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL), the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207), and the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005). DXL and TS are supported by the A*STAR Computing and Information Science (ACIS) scholarship. We thank Goh Yisheng for his contribution in the initial stage of the project. We thank members of WING and Deep Learning Lab, NUS, as well as the anonymous reviewers for the constructive feedback.
@@ Appendix B Theoretical Analysis: Reliability of $EstTrueE$
$\lim_{{FI_{C}}\to 100}{(EstTrueE)}=EstTrueE(FI_{C}=100)={TrueE}.$
@@ C.1 Multiple-choice Question (MCQ) Discussions
Table 2: MCQ results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
MMLU 32.17 30.10 1.06
BBH 41.83 22.52 93.18
Table 3: Average. estimated true accuracy results of MCQ benchmarks across models and datasets (MMLU, BBH) with two MCQ formats (Character and Textual Value). This table is the distillation result of averaging all the estimated true performance scores across models (Models) and tasks (Tasks) in Tab. 3.
@@
The FI scores across formats are illustrated in Tab. 4. There is a notable difference between the scores for character-based and textual value-based formats. Among the models, ChatGPT follows the instructions best with FI score $52.42\%$. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
@@
We investigate the impact of Chain-of-Thought (CoT) Wei et al. (2022) on mitigating format bias. The average of estimated True Accuracy scores across models and tasks for the ZS prompting and ZS-CoT prompting are shown in Tab. 5. ZS prompting achieves a higher score in the character category (40.90%) compared to ZS-CoT prompting (33.10%). Similarly, for textual value format, ZS prompting scores higher than ZS-CoT prompting. However, the $BiasF_{o}$ is lower for the ZS-CoT model ($17.42\%^{2}$) compared to the ZS model ($42.42\%^{2}$), indicating that CoT slightly decreases the format bias.
@@
Zero-shot 40.90 27.88 42.42
Zero-shot Chain-of-Thought 33.10 24.75 17.42
@@
From Tab. 2, we see that $21/32$ of the estimated $EstTrue$ results are reliable. The reliability of results in the MCQ output format varies across different models. Gemma-7B-it and Mistral-7B-it show significant unreliability in textual value format, evidenced by numerous red-marked scores due to models not following the format instructions to output correct formats. In contrast, ChatGPT’s results are significantly more reliable in the MMLU and BBH benchmarks ($7/8$), with only one unreliable result in the BBH textual format output.